OpenStudy (anonymous):

Find the slope of the line passing through the points (-7,2) and (9,6) .

OpenStudy (mertsj):

Use the slope formula.

OpenStudy (mertsj):

\[m=\frac{y _{2}-y _{1}}{x _{2}-x _{1}}\]

OpenStudy (texaschic101):

use the slope formula... slope(m) = (y2 - y1) / (x2 - x1) (-7,2) x1 = -7, y1 = 2 (9,6) x2 = 9, y2 = 6 Now sub all your stuff in and you will have the slope.

OpenStudy (anonymous):

Okay Thanks Everyone (:

OpenStudy (texaschic101):

anytime :)

OpenStudy (anonymous):

use the slope -point formula: Y-Y1=M(X-X1) this implies M(slope)=Y2-Y1/X2-X1=6-2/9+7=4/16=1/4 THEREFORE m=1/4=0.25.